OpenStudy (anonymous):

Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7.

OpenStudy (here_to_help15):

so the average rate of change between two points of a function is the slope between the two points. Remember slope is: ΔyΔx (change in y over change in x) which can be written as: y2−y1x2−x1 in this case: x2=7 and x1=4 To find the y-coordinates plug in the x values into f(x) so: y2=2(7−9)2+1 y1=2(4−9)2+1 Now just simplify those and plug them into the slope formula to find the average rate of change.

OpenStudy (here_to_help15):

the +1 should be +17

OpenStudy (anonymous):

ok thank you so much. did you mean the +1 should be 7?

OpenStudy (here_to_help15):

yesh : P

OpenStudy (anonymous):

oh ok thank you. you where a big help and i noticed you changed your picture (;

OpenStudy (here_to_help15):

Yes thanks for noticing : )